if log{(a+b)/2}=1/2(loga + log b)
prove that a=b
please provide me with a solution
 

Grade:9

1 Answers

Arun
6 years ago
log{(a+b)/2}=1/2 log(ab)
=log(ab)1/2
(a+b)/2= (ab)1/2
a+b=2(ab)1/2
(va)2 +(vb )2-2vavb=0
 
(va -vb)2=0
 
va-vb=0
 
va=vb
 
a=b